# When and How Does CLIP Enable Domain and Compositional Generalization?

### v1

@@ 6.1 The Role of Visual Embeddings
Table 3: Domain diversity increases feature sharing in the embeddings. We report the percentage point increase in top-k shared SAE features, averaged over $c\in C_{2}$ and $k\in\{5,10,15,20\}$, when comparing low diversity (Natural-only) to high diversity (Leave-out-domain, CG high-diversity). Note that we used the CLIP models using CC12M as base dataset, since we found that SAEs extracted poor features for the ImageNet-Captions models.
@@
After extracting interpretable features, we computed the percentage overlap of the $k$ most important features between pairs of domains per class. Table 3 confirms that a CLIP model that generalizes better shares indeed more features in its visual embeddings. {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = seabornblue!10, roundcorner = 4pt, ] Finding 3: CLIP shares more features in its embeddings as generalization improves.
@@ 6.2 Why Does CLIP Sometimes Fail To Generalize?
Table 4: Domain-specific captions do not explain CLIP’s poor generalization performance to unseen quickdraw classes in the CG high-diversity setting. As we replace all domain-specific captions (second row), we find that visual embeddings are more aligned (Figure 5(b)) but this does not lead to improved generalization performance.
@@
We initially hypothesized that using domain-specific captions–our training included both domain-invariant and domain-specific captions (Section 4)–might have inadvertently led CLIP to prioritize uniformity over alignment in its loss function. We suspect this occurs due to the stark visual differences between quickdraw images and those from other domains, making alignment challenging. In order to minimize the total loss in spite of this, CLIP may adopt a shortcut: prioritizing uniformity, aligning quickdraw image embeddings with the text embeddings of the domain-specific captions, while sacrificing alignment with domain-invariant captions. Consequently, we would expect a clear separation between quickdraw image embeddings and those from other domains–and indeed, we observe this (Figure 5(a)). However, this separation may come at a cost, limiting CLIP’s ability to generalize across the quickdraw domain, particularly to unseen classes.
@@
As a remedy, we trained CLIP using only domain-invariant captions. Although the visual embeddings are now better aligned (Figure 5(b)), generalization performance remains poor (second row of Table 4). Thus, the domain invariance or specificity of captions cannot solely explain CLIP’s poor performance on quickdraw images of unseen classes.
The role of shared intermediate features and circuitry
@@
For the representational analysis, we measured the representational similarity of classes across domains. Specifically, we used center kernel alignment (CKA) (Kornblith et al., 2019) with the unbiased Hilbert-Schmidt Independence Criterion (HSIC) estimator (Song et al., 2012); refer to Section C.2 for technical details.
For the mechanistic analysis, we identified the circuit for each class of each domain in the first step. Specifically, we identified the $k$ most important model components–i.e., axis-aligned neurons–by computing their indirect effect (Pearl, 2001) on the model’s predictions, following recent work (Vig et al., 2020; Schrodi et al., 2022; Meng et al., 2022; Marks et al., 2024). Intuitively, the indirect effect quantifies how much a neuron contributes to the model’s prediction. These identified neurons serve as the nodes of the graph representing the circuit. Next, we determined the $k^{\prime}$ most influential predecessors of each neuron, forming the edges of the circuits, that intuitively represent the information flow. To do this, we computed the indirect effects of the predecessors on each neuron. Refer to Section C.3 for further details on this procedure.
In the second step, we measured circuit similarity using graph similarity measures. Specifically, we measured the similarity of the circuits of each class across different domains, e.g., the circuit for quickdraw dog images vs. those for clipart, infograph, etc. We computed the layer-wise Jaccard index of the $k$ most important neurons, i.e., the nodes of the circuit. Beyond this simple node overlap measure, we used the normalized Weisfeiler-Lehman subtree graph kernel (Shervashidze et al., 2011) to capture more complex structural and hierarchical similarities. Refer to Section C.3 for details on these similarity measures.
@@
Figure 6: CLIP separates quickdraw images from other domains in its intermediate representations and circuitry, even though final visual embeddings can be aligned. We used the CLIP model trained on only domain-invariant captions from Figures 5(b) and 4 (second row). Scores are averaged over classes and higher scores mean higher representational similarity (6(a)), more shared neurons (6(b)), or more similar circuitry (6(c)).
Figure 6 confirms that the quickdraw domain noticeably differs from the other domains, both in terms of representational and mechanistic similarity, supporting our hypothesis that (compositional) generalization requires sufficient sharing of intermediate features and circuitry. {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = seabornblue!10, roundcorner = 4pt, ] Finding 4: Sufficient sharing of intermediate features and circuitry is crucial for generalization to succeed.
@@ 7 Discussion
Our results also allow us to reinterpret the main experiment from Mayilvahanan et al. (2024a, Table 1). In their experiment, the maximally achievable performance corresponds to the unaltered LAION-200M dataset, while the pruned versions of LAION-200M–where, e.g., (near) duplicates of all ImageNet-Sketch classes were removed–can be interpreted analogous to the curated compositional generalization setting with high diversity from above. They observed only a slight deterioration in performance despite these exclusions. Our result suggests that compositional generalization could be the key contributor of the sustained performance observed here, though further analysis is required to exclude other (unknown) factors.
@@
Recent works have concluded that CLIP’s OOD generalization is largely driven by its vast and diverse training distribution (Fang et al., 2022; Mayilvahanan et al., 2024b). Our experiments reaffirm that while CLIP can weakly generalize (Figure 2), a gap persists when compared to a model that has been exposed to class-specific samples of the test domain(s) (Table 1). We believe it is worth studying how narrow this gap can become and if this gap can ever be closed without access to (near) duplicate samples during training.
@@
At the time of writing, to the best of our knowledge, no large-scale domain datasets comparable to DomainNet currently exists. To address this, future work could filter a diverse dataset, such as LAION (Schuhmann et al., 2021), by domains. For example, the data filtering approach of Mayilvahanan et al. (2024b), which filters images into natural and non-natural, could be extended to various visual domains. The existing filtered natural-only subset of LAION-200M from Mayilvahanan et al. could be directly adopted as our base dataset $D_{0}$.
@@
We conducted our experiments mainly on NVIDIA RTX 2080 GPUs and estimated the total GPU hours to be approximately 25 0002500025\,00025 000. With a carbon efficiency of 0.482 kgCO2eq/kWh [^3] Estimated carbon efficiency of the global power grid in 2023. Numbers for 2024 were not released at time of writing., total emissions are estimated to be about 2 60026002\,6002 600 kgCO2eq, using the Machine Learning Impact calculator (Lacoste et al., 2019).
@@ 8 Conclusion
In this work, we analyzed when CLIP generalizes to unseen domains and domain-class pairs using systematically created training data setups (see Figure 1), identifying domain diversity as a prerequisite for both domain and compositional generalization. We showed that compositional generalization can fail in certain scenarios, i.e., even perform worse than domain generalization. We supported these findings with in-depth experiments and mechanistic analyses, offering insights into the internal workings that make generalization succeed or fail.
@@ Acknowledgements
Simon would like to thank Evgenia Rusak for an insightful discussion that sparked this project.
@@ A.1 DomainNet Examples
Figure 7 visualizes four examples from all six image domains (clipart, infograph, painting, quickdraw, real, sketch) from DomainNet (Peng et al., 2019). Some domains are more visually similar than others. For example, sketches and quickdraw are typically gray images, while paintings often contain a similar level of image detail as real (natural) images.
@@ A.3 Further Details on the Training Data Construction
We created our training datasets based on a base dataset $D_{0}$ (i.e., ImageNet-Captions, CC3M, or CC12M) which provides a large collection of (mostly) natural images along with corresponding language descriptions. To address issues related to class shifts, we then augmented the base dataset with samples from DomainNet-Real, which consists of natural images. Finally, we created different domain mixtures (Figure 1) by incorporating subsets of samples from various non-natural domains of DomainNet $D_{r}$ with $r\in\{\text{Clipart},\;\text{Infograph},\;\text{Painting},\;\text{Quickdraw},\;\text{Sketch}\}$ into the training data, as outlined in Section 3.
@@
To ensure fair performance comparisons between the different domain mixtures (Figure 1), we applied subsampling to maintain comparable final dataset sizes when adding additional domains to the training data. We designed our subsampling method to preserve the original data distribution as much as possible. Specifically, if domain $D_{i}$ contained twice as many samples as domain $D_{j}$ before subsampling, this ratio remained approximately constant afterward. Likewise, the class distribution within each domain was preserved as much as possible.
Note that the non-natural domains in DomainNet vary significantly in size, e.g., quickdraw has more than three times as many samples as clipart. Thus, we chose to only keep dataset sizes fixed within the same test domain. That is, for a given test domain, CG low-diversity and CG high-diversity datasets have the same size. However, dataset sizes may differ across test domains (i.e., CG low-diversity settings for different test domains are not necessarily of equal size). Since the Natural-only lower bound is independent of the choice of the test domain, it contains slightly fewer samples than the other mixtures.
@@ A.4 Choice of Classes $C_{2}$
We carefully chose the subset $C_{2}$ in a way that the classes are diverse and not biased towards any spurious features (e.g., color). We considered the 147 classes, with a one-to-one match in ImageNet (see Section A.7), as the possible candidates for $C_{2}$. We selected about 10% of these candidates, i.e., 15 classes. For the selection process, we randomly sampled from the set of candidates. To ensure that we adequately covered the different super-categories of DomainNet (e.g., furniture, mammal, tool, etc.), we kept only the first random sample for each category, rejecting further samples from the same category. We also manually rejected some samples if we considered them to be too similar to our existing selection. Our final selection of classes is
@@
and $C_{1}$ are all other 330 classes of DomainNet.
@@ A.5 DomainNet Evaluation Prompts
Following Radford et al. (2021), we created zero-shot weights for all 345 DomainNet Classes from these templates by taking the class-wise average over the text embeddings of all templates (marginalization to obtain the “true” object embedding) and normalizing afterwards. Formally, let $c\in\{c_{1},\dots,c_{n}\}$ be a class, $T$ be the set of templates, $t_{c}$ a template with the name of class $c$ inserted, and $g$ be the text encoder of our CLIP model. Assuming that $g$ produces $L_{2}$-normalized embeddings, we computed the zero-shot weights of $c$ as:
$\mathbf{w}_{c}=\frac{\frac{1}{|T|}\sum_{t\in T}g(t_{c})}{\left\|\frac{1}{|T|}\sum_{t\in T}g(t_{c})\right\|_{2}}.$ (3)
@@ A.7 Supervised Classifier Details
Training Details
Similar to Fang et al. (2022), we trained the supervised ResNet-50 classifiers (He et al., 2016) for 90 epochs with a batch size of 256 using SGD with Nesterov momentum, weight decay of 1e-4, momentum of 0.9, initial learning rate of 0.01 (they used 0.1) with step-wise decay by 0.1 at epochs 30, 50, and 70. We used the same image augmentations as for our CLIP models.
Joining the Class Distributions of ImageNet and DomainNet
Our final mapping assigned 450 ImageNet classes to DomainNet, including 147 one-to-one mappings. Using this mapping, we merged the class distributions of ImageNet and DomainNet. Specifically, we relabeled the 450 mapped ImageNet classes with their corresponding DomainNet labels. The remaining 550 unmatched ImageNet classes were combined with the 345 DomainNet classes, resulting in a total of 895 classes. We then trained our supervised classifiers on these 895 classes.

### v2

@@ 6.1 The Role of Visual Embeddings
Table 3: Domain diversity increases feature sharing in the embeddings. We report the percentage point increase in top-k shared SAE features, averaged over classes $c\in C_{2}$ and $k\in\{5,10,15,20\}$ (see Section C.1 for details on the computation of shared features), when comparing low diversity (Natural-only) to high diversity (Leave-out-domain, CG high-diversity). Note that we used the CLIP models using CC12M as base dataset, since we found that SAEs extracted poor features for the models trained on ImageNet-Captions.
@@
After extracting interpretable SAE features, we measured the percentage overlap of the top-$k$ most important SAE features between domain pairs for each class $c\in C_{2}$. Specifically, to identify important SAE features, we counted how often each SAE feature appeared among the top-20 highest-activating SAE features across domain-class combinations. For each class, we then calculated the percentage overlap of the top-$k$ (with $k\in\{5,10,15,20\}$) of these top-20 most frequent SAE features between the test domain and each of the other domains. These percentage overlaps were finally averaged over all values of $k$, all classes in $C_{2}$, and all pairs that include the test domain. Further details on this computation and pseudocode are provided in Section C.1.
As shown in Table 3, CLIP models trained on more diverse data—which generalize better—exhibit more cross-domain feature sharing in their visual embeddings. {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = seabornblue!10, roundcorner = 4pt, ] Finding 3: CLIP shares more features across domains in its embeddings as generalization improves.
@@ 6.2 Why Does CLIP Sometimes Fail To Generalize?
Table 4: Domain-specific captions do not explain CLIP’s poor generalization performance to unseen Quickdraw classes in the CG high-diversity setting. When all domain-specific captions are replaced (second row), we observe increased alignment in the visual embeddings (Figure 5(b)); yet, this does not translate into improved generalization performance, as measured by balanced top-1 accuracy.
@@
We initially hypothesized that using domain-specific captions—our training included both domain-invariant and domain-specific captions (Section 4)—might have inadvertently led CLIP to prioritize uniformity over alignment in its loss function. We suspect this occurs due to the stark visual differences between Quickdraw images and those from other domains, making alignment challenging. In order to minimize the total loss in spite of this, CLIP may adopt a shortcut: prioritizing uniformity, aligning Quickdraw image embeddings with the text embeddings of the domain-specific captions, while sacrificing alignment with domain-invariant captions. Consequently, we would expect a clear separation between Quickdraw image embeddings and those from other domains—and indeed, we observe this (Figure 5(a)). However, this separation may come at a cost, limiting CLIP’s ability to generalize across the Quickdraw domain, particularly to unseen classes.
@@
As a remedy, we trained CLIP using only domain-invariant captions. Although the visual embeddings are now better aligned (Figure 5(b)), generalization performance remains poor (second row of Table 4). Thus, the domain invariance or specificity of captions cannot solely explain CLIP’s poor generalization performance on unseen classes from Quickdraw.
The role of shared intermediate features and circuits
@@
Representational similarity analysis
To assess whether CLIP learns similar intermediate representations for the same class across domains, we measured representational similarity using Center Kernel Alignment (CKA) (Kornblith et al., 2019) with the unbiased Hilbert-Schmidt Independence Criterion (HSIC) estimator (Song et al., 2012); refer to Section C.2 for technical details.
Mechanistic similarity analysis
While representational similarity captures alignment at the representational level, it does not reveal whether similar computations are implemented internally. To address this, we introduce a measure of mechanistic similarity, which quantifies the similarity of the underlying circuits, i.e., the sets of model components and their interactions responsible for a specific prediction. Our analysis proceeds in two steps:
  1. 1.
    Circuit discovery: For each class-domain pair, we extracted a circuit represented as a directed acyclic graph. We first identified the top-$k$ model components—specifically, axis-aligned neurons—based on their indirect effect [^3] Intuitively, the indirect effect quantifies how much each component contributes to the model’s prediction. (Pearl, 2001) on the model’s predictions, following prior work (Vig et al., 2020; Schrodi et al., 2022; Meng et al., 2022; Marks et al., 2025). These components serve as the nodes of the circuit. Next, for each component, we identified its $k^{\prime}$ most influential predecessors (also based on indirect effect), which define the edges of the circuit. Refer to Section C.3 for further details on circuit discovery.
  2. 2.
    Circuit similarity: We then compared circuits across domains for each class—for example, comparing the circuit for “dog” of Quickdraw images vs. those of Clipart, Infograph, Painting, Real (natural images), and Sketch images. To quantify similarity, we computed the layer-wise Jaccard index of the top-$k$ components, capturing node overlap. To assess deeper structural similarity, we applied the normalized Weisfeiler-Lehman subtree graph kernel (Shervashidze et al., 2011), which captures more complex structural and hierarchical similarities. More details are provided in Section C.3.
@@
Figure 6: CLIP’s intermediate representations and circuits show a significant difference between Quickdraw and the other domains, even though final visual embeddings can be aligned. We used the CLIP model trained on only domain-invariant captions from Figures 5(b) and 4 (second row). Scores are averaged over classes and higher scores mean higher representational similarity (6(a)), more shared neurons (6(b)), or more similar circuits (6(c)).
Figure 6 confirms that the Quickdraw domain noticeably differs from the other domains, both in terms of representational and mechanistic similarity, supporting our hypothesis that (compositional) generalization requires sufficient sharing of intermediate features and circuits. {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = seabornblue!10, roundcorner = 4pt, ] Finding 4: Sufficient sharing of intermediate features and circuits is crucial for generalization to succeed.
@@ 7 Discussion
Our results also allow us to reinterpret the main experiment from Mayilvahanan et al. (2024, Table 1). In their experiment, the maximally achievable performance corresponds to the unaltered LAION-200M dataset, while the pruned versions of LAION-200M—where, e.g., (near) duplicates of all ImageNet-Sketch classes were removed—can be interpreted analogous to the curated compositional generalization setting with high diversity from above. They observed only a slight deterioration in performance despite these exclusions. Our result suggests that compositional generalization could be the key contributor of the sustained performance observed here, though further analysis is required to exclude other (unknown) factors.
@@
Recent studies have concluded that CLIP’s OOD generalization is largely driven by its vast and diverse training distribution (Fang et al., 2022; Mayilvahanan et al., 2025). Our experiments reaffirm this: While CLIP can weakly generalize (Figure 2), a generalization gap persists when compared to models that have seen class samples of the test domain(s) during training (Table 1). We believe it is worth studying how narrow this generalization gap can become and if it can ever be closed without access to (near) duplicate samples during training—both from empirical and theoretical perspectives.
@@
In this work, we disentangled the effects of dataset size and domain diversity—two factors that are typically confounded—and showed that domain diversity is a key driver of CLIP’s generalization. However, dataset size alone may also impact CLIP’s generalization performance.
At the time of writing, to the best of our knowledge, no large-scale domain datasets comparable to the size and diversity of DomainNet exists. To address this, future work could filter a large, diverse dataset, such as LAION (Schuhmann et al., 2021), by domains. For example, the data filtering approach proposed by Mayilvahanan et al. (2025), which classifies images into natural and non-natural, could be extended to classify various visual domains. Note that the existing filtered natural-only subset of LAION-200M curated by Mayilvahanan et al. could be directly adopted as the base dataset $D_{0}$.
@@
Dataset quality
Besides diversity, the quality of the datasets used for training also likely plays a crucial role in achieving robust generalization. For example, Nguyen et al. (2022) showed that combining multiple datasets does not necessarily result in improved robustness; in fact, the robustness of the best-performing dataset can even be diluted by lower-quality datasets. To date, dataset quality remains only loosely addressed in much of the current literature, largely due to the difficulty to quantify it. Thus, we leave further investigation of this aspect for future work.
@@
We conducted our experiments mainly on NVIDIA RTX 2080 GPUs and estimated the total GPU hours to be approximately 25 0002500025\,00025 000. With a carbon efficiency of 0.321 kgCO2eq/kWh [^4] Estimated carbon efficiency of the German power grid in 2024 according to https://www.nowtricity.com/country/germany/, total emissions are estimated to be about 1 72517251\,7251 725 kgCO2eq, using the Machine Learning Impact calculator (Lacoste et al., 2019).
@@ 8 Conclusion
In this work, we analyzed when CLIP generalizes to unseen domains and subsets thereof using systematically created training data setups (see Figure 1), identifying domain diversity as a prerequisite for both domain and compositional generalization. We showed that compositional generalization can fail in certain scenarios, i.e., even perform worse than domain generalization. We supported these findings with in-depth data-centric experiments and mechanistic analyses, offering insights into the internal workings that make generalization succeed or fail.
@@ Acknowledgements
We would like to thank Evgenia Rusak, Thaddäus Wiedemer, and Prasanna Mayilvahanan for an insightful discussion.
@@ A.1 DomainNet Examples
Figure 7 visualizes four examples from all six image domains (Clipart, Infograph, Painting, Quickdraw, Real, Sketch) from DomainNet (Peng et al., 2019). Some domains are more visually similar than others. For example, sketches and Quickdraw images are typically gray, while paintings often contain a similar level of image detail similar to that of Real (natural) images.
@@ A.3 Further Details on the Training Data Construction
We created our training datasets based on a base dataset $D_{0}$ (i.e., ImageNet-Captions, CC3M, or CC12M) which provides a large collection of (mostly) natural images along with corresponding language descriptions. To mitigate effects stemming from to class shifts, we augmented the base dataset with samples from DomainNet-Real, which consists of natural images. Finally, we created different domain mixtures (see Figure 1) by incorporating subsets of samples from various non-natural domains of DomainNet $D_{r}$ with $r\in\{\text{Clipart},\;\text{Infograph},\;\text{Painting},\;\text{Quickdraw},\;\text{Sketch}\}$ into the training data, as outlined in Section 3.
@@
To ensure fair performance comparisons between the different training data setups (see Figure 1), we applied subsampling to maintain comparable final dataset sizes when adding additional domains to the training data. We designed our subsampling method to preserve the original data distribution as much as possible. Specifically, if domain $D_{i}$ contained twice as many samples as domain $D_{j}$ before subsampling, this ratio remained approximately constant afterward. Likewise, the class distribution within each domain was preserved as much as possible.
Note that the non-natural domains in DomainNet vary significantly in size, e.g., Quickdraw has more than three times as many samples as Clipart. Thus, we chose to only keep dataset sizes fixed within the same test domain. That is, for a given test domain, CG low-diversity, CG high-diversity and Leave-out-domain datasets have the same size. However, dataset sizes may differ across test domains (i.e., CG low-diversity settings for different test domains are not necessarily of equal size). Since the Natural-only lower bound is independent of the choice of the test domain, it contains slightly fewer samples than the other mixtures.
Joining the Class Distributions of ImageNet and DomainNet
Our final mapping assigned 450 ImageNet classes to DomainNet, including 147 one-to-one mappings. Using this mapping, we merged the class distributions of ImageNet and DomainNet. Specifically, we relabeled the 450 mapped ImageNet classes with their corresponding DomainNet labels.
@@ A.4 Choice of Classes $C_{2}$
We carefully chose the subset $C_{2}$ in a way that the classes are diverse and not biased towards any spurious features (e.g., color). We considered the 147 classes, with a one-to-one match in ImageNet (see Section A.3), as the possible candidates for $C_{2}$. We selected about 10% of these candidates, i.e., 15 classes. For the selection process, we randomly sampled from the set of candidates. To ensure that we adequately covered the different super-categories of DomainNet (e.g., furniture, mammal, tool), we kept only the first random sample for each category, rejecting further samples from the same category. We also manually rejected some samples if we considered them to be too similar to our existing selection. Our final selection of classes is
@@
and $C_{1}$ are the remaining 330 classes of DomainNet.
@@ A.5 DomainNet Evaluation Prompts
Following Radford et al. (2021), we created zero-shot weights for all 345 DomainNet Classes from these templates by taking the class-wise average over the text embeddings of all templates (marginalization to obtain the “true” object embedding) and normalizing afterwards. Formally, let $c\in\{c_{1},\dots,c_{n}\}$ be a class, $T$ be the set of templates, $t_{c}$ a template with the name of class $c$ inserted, and $g$ be the text encoder of our CLIP model. Assuming that the text encoder $g$ produces $L_{2}$-normalized embeddings, we compute the zero-shot weights of class $c$ as:
$\mathbf{w}_{c}=\frac{\frac{1}{|T|}\sum_{t\in T}g(t_{c})}{\left\|\frac{1}{|T|}\sum_{t\in T}g(t_{c})\right\|_{2}}\quad.$ (3)
@@ A.7 Supervised Classifier Details
Similar to Fang et al. (2022), we trained the supervised ResNet-50 classifiers (He et al., 2016) for 90 epochs with a batch size of 256 using SGD with Nesterov momentum, weight decay of 1e-4, momentum of 0.9, initial learning rate of 0.01 (they used 0.1) with step-wise decay by 0.1 at epochs 30, 50, and 70. We used the same image augmentations as for our CLIP models. We trained our supervised classifiers on the 895 classes—550 unmatched ImageNet classes combined with the 345 DomainNet classes—see Section A.3 for details